# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v1

@@ 4.2.2. Datasets
Niklaus et al. (2021) introduce a multilingual LJP dataset from the Federal Supreme Court of Switzerland (FSCS), containing over 85,000 cases in German, French, and Italian. The dataset is annotated with publication years, legal areas, and cantons of origin, making it suitable for NLP applications in judgment prediction.
@@
Semo et al. (2022) introduce the first LJP dataset focused on class action cases in the United States. The dataset targets predicting outcomes based on plaintiffs’ complaints rather than court-written fact summaries, involving a rule-based extraction system to identify relevant text spans from complaints.
Almuslim and Inkpen (2022) construct a dataset for LJP within the Saskatchewan Court of Appeal. They collected and labeled 3,670 documents with case outcomes (‘allow’ or ‘dismiss’) using a two-step pattern matching and keyword-based validation, ensuring label accuracy through manual annotation. This dataset supports research in the Canadian legal system and aids in developing predictive models for legal judgments.
4.3. Legal Text Classification
4.3.1. Task
@@
LTC is an important task within the domain of NLP that involves categorizing legal documents based on their content, a foundational aspect of building intelligent legal systems. With the exponential growth of legal documents, it has become increasingly challenging for legal professionals to efficiently locate relevant rulings in similar cases for argumentation. LTC addresses this challenge by automatically associating legal texts with predefined categories, such as criminal, civil, or administrative cases, thereby simplifying legal research and decision-making processes.
@@
In the legal field, this process is often referred to as predictive coding, where ML algorithms are trained through supervised learning to classify documents into specific categories. The broader task of text classification in NLP involves assigning one or multiple categories to a document from a set of predefined options, and it can take various forms, including binary classification, multi-class classification, and multi-label classification. Legal document classification often falls under large multi-label text classification, where the label space can consist of thousands of potential categories, adding significant complexity to the task (Shaheen et al., 2020). This subsection explores the methodologies and advancements in this area.
@@
DL typically requires extensive data to yield effective results, but MTL offers a potential solution to the data scarcity problem. Elnaggar et al. (2018) leverage transfer learning and MTL to perform tasks like translation and multi-label classification within legal document corpora. They utilize the MultiModel algorithm (Kaiser et al., 2017), which uses a fully convolutional sequence-to-sequence architecture integrating different modality nets. Their model processes legal texts through a unified embedding, enabling efficient task switching and promoting generalization across different legal tasks, effectively tackling data scarcity in the legal field.
@@
Wei et al. (2018) investigate the application of CNNs for text classification in legal document review, comparing their performance with SVMs on four datasets from real legal cases. They found that CNNs perform better with larger training datasets and offer a more stable growth trend compared to SVMs. The study highlighted challenges such as sequence length limitations and the need for explainability in predictive analysis, suggesting improvements like integrating annotated sentences with full text to enhance sentence relevance identification.
@@
Lee and Lee (2019) focuse on legal document classification in the Korean language and compare three different DL approaches: CNN with ASCII encoding, CNN with Word2Vec embeddings, and RNN with Word2Vec embeddings. The classification models are used to classify case data into civil, criminal, and administrative. Using a dataset of nearly 60,000 past case documents, the study finds that the RNN model with Word2Vec embedding achieves the highest classification accuracy.
@@
Bambroo and Awasthi (2021) introduce an architecture that integrates long attention mechanisms with a distilled BERT model pre-trained on legal domain-specific corpora. Their model employs a combination of local windowed attention and task-motivated global attention to handle inputs up to eight times longer than standard BERT models. The architecture, based on DistilBERT (Sanh et al., 2020) and incorporating LongformerSelf-Attention, is optimized for legal document classification, outperforming fine-tuned BERT and other transformer-based models in both speed and effectiveness.
@@
Fragkogiannis et al. (2023) propose a method to improve classification of pages within lengthy documents by leveraging sequential context from preceding pages. They enhance the input to pre-trained models like BERT (Devlin et al., 2019) by appending special tokens representing the predicted page type of the previous page, enabling more context-aware classification without modifying the model architecture. Experiments on legal datasets demonstrate improvements compared to non-recurrent setups.
@@
Mamooler et al. (2022) propose an active learning pipeline for fine-tuning PLMs for LTC. They address challenges of specialized vocabulary and high annotation costs. Their method involves continued pre-training of RoBERTa (Liu et al., 2019) on legal texts, knowledge distillation using a pre-trained sentence transformer, and an efficient initial sampling strategy by clustering unlabeled data. This approach reduces the number of labeling actions required and improves efficiency in adapting models to LTC tasks.
4.3.2. Datasets
LTC datasets are characterized by their domain-specific vocabulary and multi-label nature, requiring models to interpret complex legal texts and categorize them into single or multiple legal themes.
@@ 4.3.2. Datasets
Tuggener et al. (2020) introduce LEDGAR, a multi-label corpus of legal provisions from contracts scraped from the U.S. Securities and Exchange Commission’s website. The dataset includes over 846,000 provisions across 60,540 contracts, with an extensive label set suitable for text classification and legal studies, aiding in developing advanced legal NLP models.
Chalkidis et al. (2021) present MULTI-EURLEX, a multilingual dataset containing 65,000 EU laws translated into 23 official EU languages, annotated with EUROVOC labels. The dataset emphasizes temporal concept drift by adopting chronological splits, enhancing its utility for sophisticated LTC tasks requiring understanding nuanced legal terms across different time periods.
Papaloukas et al. (2021) introduce the Greek Legal Code dataset, categorizing approximately 47,000 Greek legislative documents into a detailed multi-level classification system. The dataset is structured into volumes, chapters, and subjects, each containing diverse legal documents from Greek legislation history, supporting LTC in the Greek legal domain.
@@
Graham et al. (2023) develop a domain-specific dataset for LTC focusing on deontic modalities in contract sentences. They manually annotated contract sentences to train models for identifying deontic sentences like permissions, obligations, and prohibitions. The corpus, derived from the Contract Understanding Atticus Dataset (CUAD) (Hendrycks et al., 2021), provides a resource for studying functional categories crucial for legal analysis.
4.4. Legal Document Sumarrization
4.4.1. Task
@@
LDS is a specialized branch of automatic summarization that focuses on condensing legal texts, such as court judgments, into clear and informative summaries. Unlike general text summarization, which extracts key details without following specific formatting rules, LDS must account for the distinct structure and specialized content of legal documents. These documents often include complex details like article numbers, statutory language, and citations that are critical for presenting the legal arguments and decisions accurately. The natural complexity of legal texts, characterized by their extensive length and detailed internal structures such as sections, articles, and paragraphs in statutes—demands tailored summarization techniques. This need is further emphasized by the hierarchical importance of documents based on their judicial origin, where the interpretation of texts can vary significantly between higher and lower court opinions (Kanapala et al., 2019).
@@
LDS can be approached through extractive and abstractive methods. Extractive summarization techniques in LDS focus on identifying and extracting the most critical sentences or phrases directly from the text, maintaining the original wording and meaning. In contrast, abstractive summarization involves generating new sentences that paraphrase the most important information, aiming for conciseness and coherence while ensuring that the essence of the legal text is preserved. This subsection will explore existing approaches to LDS and show the small differences that set it apart from more general summarization methods and discussing the challenges and solutions specific to the legal domain.
@@
Several systems have been specifically designed to summarize legal documents. One of the first systems in this field was the Fast Legal EXpert CONsultant (FLEXICON), created by Gelbart and Smith in 1991 (Gelbart and Smith, 1991a). FLEXICON utilizes a keyword-based approach (Gelbart and Smith, 1991b), scanning a comprehensive database of terms to pinpoint crucial segments of text. Following this, Moens et al. (1999) introduced the SALOMON system in 1999, which employs cosine similarity to cluster similar text regions, aiming to highlight relevant topics within the documents. This method aligns with other abstraction-oriented techniques seen in the work of Erkan and Radev (2004). Another system, LetSum, devised by Farzindar and Lapalme in 2004 (Farzindar, 2004), also adopts a keyword-centric strategy but uses “cue phrases” to identify text related to specific themes such as ‘Introduction’, ‘Context’, and ‘Conclusion’. Although LetSum was fairly successful in mimicking human-generated summaries, it tended to produce summaries that were excessively lengthy.
@@
Zhong et al. (2019) propose an automatic extractive summarization system for legal cases concerning Post-traumatic Stress Disorder from the US Board of Veterans’ Appeals. It employs a train-attribute-mask pipeline using a CNN classifier to iteratively select predictive sentences from case texts.
@@
Nguyen et al. (2021) propose an RL framework to enhance deep summarization models for the legal domain, utilizing Proximal Policy Optimization with a reward function that integrates both lexical and semantic criteria. They fine-tune an extractive summarization backbone based on BERTSUM (Liu, 2019), employing a reward model that includes lexical, sentence, and keyword-level semantics to produce better legal summaries. Schraagen et al. (2022) apply an RL approach with a Bi-LSTM and a deep learning approach based on the BART transformer model to abstractive summarization of the Dutch case verdict database Rechtspraak.nl, combining extractive and abstractive summarization to retain core facts while creating concise summaries.
@@
Zhong and Litman (2022) focus on extractive summarization of legal case decisions, proposing an unsupervised graph-based ranking model that leverages a reweighting algorithm to utilize document structure properties. They introduce a reweighting algorithm to improve sentence selection in the HipoRank model (Dong et al., 2021). It aims to reduce redundancy and enhance the selection of argumentative sentences from underrepresented sections.
@@
Moro et al. (2023) introduce a transfer learning approach that combines extractive and abstractive summarization techniques to address the lack of labeled legal summarization datasets, outperforming previous results on the Australian Legal Case Reports dataset and establishing a new baseline for abstractive summarization.
@@
Jain et al. (2024) propose a sentence scoring approach, DCESumm, which combines supervised sentence-level summary relevance prediction with unsupervised clustering-based document-level score enhancement. They utilize a Legal BERT-based Multi-Layer Perceptron (MLP) model to predict the summary relevance of each sentence, refining scores through deep embedded sentence clustering to enhance the selection process by considering the global context of the document.
@@
Liu et al. (2024) present Common Law Court Judgment Summarization (CLSum), a pioneering dataset for summarizing multi-jurisdictional common law court judgments, leveraging large language models for data augmentation, summary generation, and evaluation. They employ a two-stage summarization process with techniques like sparse attention mechanisms and efficient training methods to process lengthy legal documents within limited computational resources.
4.4.2. Datasets
@@
LDS datasets are largely built from structured court proceedings and decisions, providing rich sources for both extractive and abstractive summarization methods. These datasets often emphasize abstractive summarization to achieve concise, readable summaries that transform the original legal language into more accessible forms (Shen et al., 2022).
Zhong et al. (2019) develop a dataset from 972,522 Board of Veterans’ Appeals decisions, focusing on single-issue cases related to Post-traumatic Stress Disorder. The dataset consists of 112 carefully sampled decisions, annotated by legal experts to capture key information such as ‘Issue’, ‘Procedural History’, ‘Service History’, ‘Outcome’, ‘Reasoning’, and ‘Evidential Support’.
@@
Shen et al. (2022) introduce Multi-LexSum, an abstractive summarization dataset tailored for U.S. federal civil rights lawsuits, containing 40,000 source documents and 9,000 expert-written summaries of diverse lengths, providing a rich resource for testing advanced summarization models.
@@
Liu et al. (2024) publish CLSum, a dataset designed for summarizing multi-jurisdictional common law court judgments from Australia, Hong Kong, the United Kingdom, and Canada. This dataset leverages large language models for data augmentation and incorporates legal knowledge to enhance summary generation and evaluation. This dataset addresses the challenge of sparse labeled data in legal domains. CLSum includes a comprehensive collection of judgments and summaries from prominent court websites. It employs novel techniques to enrich training sets and improve model performance in few-shot and zero-shot learning scenarios.
4.5. Legal Named Entity Recognition
4.5.1. Task
@@
NER is a fundamental task in NLP that involves identifying specific segments of text and classifying them into predefined categories such as ‘organization’, ‘person’, and ‘location’ (Layegh et al., 2023). In the legal domain, NER extends to specialized recognition tasks that focus on extracting entities unique to legal texts, such as laws, legal norms, and procedural terms. This specialized form of NER is crucial for structuring legal documents and enhancing legal IR systems. Unlike general NER systems that handle common entity types, legal NER must navigate the complex language and structured format of legal documents, underscoring the need for systems and methodologies specifically tailored to the legal context.

### v2

@@ 6.2. Datasets
Niklaus et al. (2021) introduce a multilingual LJP dataset from the Federal Supreme Court of Switzerland (FSCS), containing over 85,000 cases in German, French, and Italian. The dataset is annotated with publication years, legal areas, and cantons of origin, making it suitable for NLP applications in judgement prediction.
@@
Semo et al. (2022) introduce the first LJP dataset focused on class action cases in the US. The dataset targets predicting outcomes based on plaintiffs’ complaints rather than court-written fact summaries, involving a rule-based extraction system to identify relevant text spans from complaints.
7. Legal Text Classification
@@
LTC is an important task within the domain of NLP that involves categorising legal documents based on their content, a foundational aspect of building intelligent legal systems. With the exponential growth of legal documents, it has become increasingly challenging for legal professionals to locate relevant rulings in similar cases for argumentation. LTC addresses this challenge by automatically associating legal texts with predefined categories, such as criminal, civil, or administrative cases, thereby simplifying legal research and decision-making processes.
@@
In the legal field, this process is often referred to as predictive coding, where ML algorithms are trained through supervised learning to classify documents into specific categories. The broader task of text classification in NLP involves assigning one or multiple categories to a document from a set of predefined options, and it can take various forms, including binary classification, multi-class classification, and multi-label classification. Legal document classification often falls under large multi-label text classification, where the label space can consist of thousands of potential categories, adding complexity to the task (Shaheen et al., 2020).
7.1. Task
@@
DL typically requires extensive data to yield effective results, but MTL offers a potential solution to the data scarcity problem. Elnaggar et al. (2018) leverage transfer learning and MTL to perform tasks, such as translation and multi-label classification within legal document corpora. They utilise the MultiModel algorithm (Kaiser et al., 2017), which uses a fully convolutional sequence-to-sequence architecture integrating different modality nets. Their model processes legal texts by mapping them into a shared embedding space, allowing task switching and supporting generalisation across legal tasks, thereby addressing data scarcity in the legal field.
@@
Wei et al. (2018) investigate the application of CNNs for text classification in legal document review, comparing their performance with SVMs on four datasets from real legal cases. They found that CNNs perform better with larger training datasets and offer a more stable growth trend compared to SVMs. The study highlighted challenges such as sequence length limitations and the need for explainability in predictive analysis, suggesting improvements, such as integrating annotated sentences with full text to enhance sentence relevance identification.
@@
Lee and Lee (2019) focus on LTC in the Korean language and compare three different DL approaches: CNN with ASCII encoding, CNN with Word2Vec embeddings, and RNN with Word2Vec embeddings. The classification models are used to classify case data into civil, criminal, and administrative. Using a dataset of nearly 60,000 past case documents, the study finds that the RNN model with Word2Vec embedding achieves the highest classification accuracy.
@@
Bambroo and Awasthi (2021) introduce an architecture that integrates long attention mechanisms with a distilled BERT model pre-trained on legal domain-specific corpora. Their model employs a combination of local windowed attention and task-motivated global attention to handle inputs up to eight times longer than standard BERT models. The architecture, based on DistilBERT (Sanh et al., 2020) and incorporating LongformerSelf-Attention, is optimised for legal document classification, outperforming fine-tuned BERT and other transformer-based models in both speed and performance.
@@
Mamooler et al. (2022) propose an active learning pipeline for fine-tuning PLMs for LTC. They address challenges of specialised vocabulary and high annotation costs. Their method involves continued pre-training of RoBERTa on legal texts, knowledge distillation using a pre-trained sentence transformer, and an initial sampling strategy by clustering unlabelled data. This approach reduces the number of labelling actions required for LTC tasks.
Grabmair et al. (2015) introduce LUIMA, a system designed for conceptual legal document retrieval, focusing on vaccine injury decisions. LUIMA employs a multi-level classification pipeline: rule-based sub-sentence annotations tag legal concepts, such as terms and mentions, while machine learning classifiers categorise sentences into argumentative roles, such as legal rules or evidence-based findings. These annotations feed into a sentence-level indexing process using Apache Lucene, enabling semantic querying beyond traditional keyword search. A learning-to-rank module refines the retrieved results by leveraging features, such as sentence match counts and similarity scores.
Galassi et al. (2024) explore the binary classification task of detecting potentially unfair clauses in online Terms of Service (ToS) at the sentence level. Sentences are labelled as either unfair/potentially unfair or fair. Starting from an English-trained model, the authors compare four cross-lingual approaches: (1) training separate models on each language, (2) projecting annotations from English to another language, (3) translating English training documents and using their original annotations, and (4) translating query documents into English during prediction. The evaluation shows scenario (4) achieves similar or better results than scenario (1) when translation quality is high. If the translation quality is low, scenario (1) can produce better results. Scenarios (2) and (3) perform slightly below scenario (1) but remain viable alternatives.
Rhetorical role labelling is the task of classifying sentences in legal documents based on their functional roles, such as fact, argument, or ruling, to structure and analyse judicial decisions. The Artificial Intelligence for Legal Assistance (AILA) shared task series focuses on advancing legal NLP by introducing datasets and challenges that address core legal text processing tasks. The latest edition, AILA 2021 (Parikh et al., 2022), featured a rhetorical role labelling task that required classifying sentences into seven predefined roles. Building on these challenges, DeepRhole (Bhattacharya et al., 2023) introduces a transformer-based framework that fine-tunes domain-adapted models such as LegalBERT to capture the semantics of legal texts. The system applies different word embedding strategies, including Law2Vec and embeddings from Google News, to model legal language variability. An inter-annotator study examines the subjectivity in rhetorical role assignment, further assessing the model’s performance in different judicial contexts.
7.2. Datasets
LTC datasets are characterised by their domain-specific vocabulary and multi-label nature, requiring models to interpret complex legal texts and categorise them into single or multiple legal themes.
@@ 7.2. Datasets
Tuggener et al. (2020) introduce LEDGAR, a multi-label corpus of legal provisions from contracts scraped from the U.S. Securities and Exchange Commission’s website. The dataset includes over 846,000 provisions across 60,540 contracts, with an extensive label set suitable for text classification and legal studies, aiding in developing legal NLP models.
Chalkidis et al. (2021) present MULTI-EURLEX, a multilingual dataset containing 65,000 EU laws translated into 23 official EU languages, annotated with EuroVoc labels. The dataset emphasises temporal concept drift by adopting chronological splits, enhancing its utility for sophisticated LTC tasks requiring understanding legal terms across different time periods.
Papaloukas et al. (2021) introduce the Greek Legal Code dataset, categorising approximately 47,000 Greek legislative documents into a detailed multi-level classification system. The dataset is structured into volumes, chapters, and subjects, each containing diverse legal documents from Greek legislation history, supporting LTC in the Greek legal domain.
@@
Graham et al. (2023) develop a domain-specific dataset for LTC focusing on deontic modalities in contract sentences. They manually annotated contract sentences to train models for identifying deontic sentences, such as permissions, obligations, and prohibitions. The corpus, derived from the Contract Understanding Atticus Dataset (CUAD) (Hendrycks et al., 2021), provides a resource for studying functional categories for legal analysis.
Galassi et al. (2024) extend the Claudette [^6] https://claudette.eui.eu corpus from 25 to 50 ToS per language (English, German, Italian, Polish). Each sentence is labelled according to nine categories of potential unfairness, with annotations indicating the degree of unfairness. The dataset was carefully compiled based on language availability, structural similarity, and version correspondence. Cross-lingual analysis revealed notable differences between languages—particularly in German—including variations in length, structure, missing clauses, and legal terminology, reflecting manual drafting adjustments rather than simple automated translations.
8. Legal Document Summarisation
@@
LDS is a specialised branch of automatic summarisation that focuses on condensing legal texts, such as court judgements, into clear and informative summaries. Unlike general text summarisation, which extracts key details without following specific formatting rules, LDS must account for the distinct structure and specialised content of legal documents. These documents often include complex details, such as article numbers, statutory language, and citations that are critical for presenting the legal arguments and decisions accurately. The natural complexity of legal texts, characterised by their extensive length and detailed internal structures such as sections, articles, and paragraphs in statutes—demands tailored summarisation techniques. This need is further emphasised by the hierarchical importance of documents based on their judicial origin, where the interpretation of texts can vary between higher and lower court opinions (Kanapala et al., 2019).
@@
LDS can be approached through extractive and abstractive methods. Extractive summarisation techniques focus on identifying and extracting the most critical sentences or phrases directly from the text, maintaining the original wording and meaning. In contrast, abstractive summarisation involves generating new sentences that paraphrase the most important information, aiming for conciseness while ensuring that the essence of the legal text is preserved.
8.1. Task
@@
Several systems have been specifically designed to summarise legal documents. One of the first systems in this field was the Fast Legal EXpert CONsultant (FLEXICON), created by Gelbart and Smith in 1991 (Gelbart and Smith, 1991a). FLEXICON utilises a keyword-based approach (Gelbart and Smith, 1991b), scanning a database of terms to pinpoint crucial segments of text. Following this, Moens et al. (1999) introduced the SALOMON system in 1999, which employs cosine similarity to cluster similar text regions, aiming to highlight relevant topics within the documents. This method aligns with other abstraction-oriented techniques seen in the work of Erkan and Radev (2004). Another system, LetSum, devised by Farzindar and Lapalme in 2004 (Farzindar, 2004), also adopts a keyword-centric strategy but uses “cue phrases” to identify text related to specific themes such as ‘Introduction’, ‘Context’, and ‘Conclusion’. Although LetSum was fairly successful in mimicking human-generated summaries, it tended to produce summaries that were excessively lengthy.
@@
Nguyen et al. (2021) propose an RL framework to enhance deep summarisation models for the legal domain, utilising Proximal Policy Optimisation with a reward function that integrates both lexical and semantic criteria. They fine-tune an extractive summarisation backbone based on BERTSUM (Liu, 2019), employing a reward model that includes lexical, sentence, and keyword-level semantics to produce better legal summaries. Schraagen et al. (2022) apply an RL approach with a Bi-LSTM and a deep learning approach based on the BART transformer model to abstractive summarisation of the Dutch case verdict database Rechtspraak.nl, combining extractive and abstractive summarisation to retain core facts while creating concise summaries.
@@
Zhong and Litman (2022) focus on extractive summarisation of legal case decisions, proposing an unsupervised graph-based ranking model that leverages a reweighting algorithm to utilise document structure properties. They introduce a reweighting algorithm to improve sentence selection in the HipoRank model (Dong et al., 2021). It aims to reduce redundancy and enhance the selection of argumentative sentences from underrepresented sections.
@@
Moro et al. (2023) introduce a transfer learning approach that combines extractive and abstractive summarisation techniques to address the lack of labelled legal summarisation datasets, outperforming previous results on the Australian Legal Case Reports dataset and establishing a new baseline for abstractive summarisation.
@@
Jain et al. (2024) propose a sentence scoring approach, DCESumm, which combines supervised sentence-level summary relevance prediction with unsupervised clustering-based document-level score enhancement. It utilises a Legal BERT-based Multi-Layer Perceptron model to predict the summary relevance of each sentence, refining scores through deep embedded sentence clustering to enhance the selection process by considering the global context of the document.
@@
Liu et al. (2024) present Common Law Court Judgement Summarisation (CLSum), a pioneering dataset for summarising multi-jurisdictional common law court judgements, leveraging LLMs for data augmentation, summary generation, and evaluation. They employ a two-stage summarisation process with techniques, such as sparse attention mechanisms and efficient training methods to process lengthy legal documents within limited computational resources.
8.2. Datasets
@@
LDS datasets are largely built from structured court proceedings and decisions, providing rich sources for both extractive and abstractive summarisation methods. These datasets often use abstractive summarisation to achieve concise, readable summaries that transform the original legal language into more accessible forms (Shen et al., 2022).
@@
Shen et al. (2022) introduce Multi-LexSum, an abstractive summarisation dataset tailored for U.S. federal civil rights lawsuits, containing 40,000 source documents and 9,000 expert-written summaries of diverse lengths, providing a rich resource for testing summarisation models.
@@
Liu et al. (2024) publish CLSum, a dataset designed for summarising multi-jurisdictional common law court judgements from Australia, Hong Kong, the United Kingdom, and Canada. This dataset leverages LLMs for data augmentation and incorporates legal knowledge to enhance summary generation and evaluation. This dataset addresses the challenge of sparse labelled data in legal domains. CLSum includes a collection of judgements and summaries from prominent court websites. It employs novel techniques to enrich training sets and improve model performance in few-shot and zero-shot learning scenarios.
9. Legal Named Entity Recognition
@@
NER is a fundamental task in NLP that involves identifying specific segments of text and classifying them into predefined categories such as ‘organisation’, ‘person’, and ‘location’ (Layegh et al., 2023). In the legal domain, NER extends to specialised recognition tasks that focus on extracting entities unique to legal texts, such as laws, legal norms, and procedural terms. This specialised form of NER is crucial for structuring legal documents and enhancing legal IR systems. Unlike general NER systems that handle common entity types, legal NER must navigate the complex language and structured format of legal documents, underscoring the need for systems and methodologies specifically tailored to the legal context.
9.1. Task